OpenStudy (anonymous):

What percent of 28 is 14?

OpenStudy (anonymous):

\[ (x)\frac{ 28}{100 }=14\\ (x)\frac{ 7}{25 } = 14\\ x=50\]

OpenStudy (anonymous):

Let the percentage be x: So : \[x \% of 28 = 14 \implies \frac{x}{100} \times 28 = 14 \implies x = \frac{14 \times 100}{28} = Solve \; \; Yourself\]

OpenStudy (anonymous):

Thanks ya!